# Representative Action Selection for Large Action Space Meta-Bandits

### v2

@@ 3.1 Regret Bounds of Algorithm
$\begin{split}\mathbb{E}_{\theta,\mathcal{A}}[\operatorname*{Regret}]\geq&\left(\min_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]-C\epsilon\sqrt{\log|\mathcal{R}|}\right)\\ &\times c\cdot\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\right)^{1/2}.\end{split}$
3.2 Regret Bound for Sub-Gaussian Process
@@
Let $\gamma_{2}(\mathcal{S})$ denote Talagrand’s chaining functional (Talagrand, 2014) for a set $\mathcal{S}\subset\mathbb{R}^{n}$, $n\in\mathbb{N}$, equipped with the Euclidean norm. It is the sharpest known bound that for a sub-Gaussian process indexed by $\mathcal{S}$, there exists a constant $c>0$ such that $\mathbb{E}\left[\max_{a\in\mathcal{S}}\mu_{a}\right]\leq c\gamma_{2}(\mathcal{S})$. Using this functional, we establish a regret bound for our algorithm:
@@
Let $\{\mu_{a}\}_{a\in\mathcal{A}_{\operatorname*{full}}}$ be a mean-zero sub-Gaussian process. Consider a partition :=$\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space. Let $\mathcal{A}$ be output of Algorithm 1. Then, for a constant $C>0$,
4 Generalization and Empirical Validation
In general cases, the decision-maker may not have explicit access to the full structure of the action space, especially in high-dimensional settings. Instead, they are given a list of actions and can observe the expected outcomes of these actions through sampling. We therefore adopt a dimension-free view by treating the expected outcomes $\{\mu_{a}\}_{a\in\mathcal{A}_{\operatorname*{full}}}$ as a family of random variables indexed by an abstract set $\mathcal{A}_{\operatorname*{full}}$. In a single bandit, these expected outcomes define an outcome function $f$ over the action space $\mathcal{A}_{\operatorname*{full}}$:
$f(a):=\mu_{a},\quad\forall a\in\mathcal{A}_{\operatorname*{full}}.$
[…]
$k(a,a^{\prime})=\sum\nolimits_{i\leq\infty}\lambda_{i}\phi_{i}(a)\phi_{i}(a^{\prime}),$
where $(\phi_{i})_{i\leq\infty}$ are orthonormal eigenfunctions under $\mathbb{P}$, and $(\lambda_{i})_{i\leq\infty}$ are non-negative eigenvalues. Let outcome $f$ be of the form $f(\cdot)=\sum_{i=1}^{N}\alpha_{i}k(\cdot,a_{i})$ for some integer $N\geq 1$, and a set of points $\{a_{i}\}_{i=1}^{N}\subset\mathcal{A}_{\operatorname*{full}}$ and a weight vector $\alpha\in\mathbb{R}^{N}$. The function $f$ can be rewritten as
$\mu_{a}=f(a)=\langle\mathbf{f},\Phi(a)\rangle,$
where $\mathbf{f}$ is a vector of coefficients, and $\Phi(a)$ usually refereed as a feature map, has entries $\Phi_{i}(a)=\sqrt{\lambda_{i}}\phi_{i}(a)$. If the coefficients $\{\mathbf{f}_{i}\}_{i=1}^{\infty}$ are i.i.d. Gaussian, this formulation can be represented as the canonical Gaussian process model in Equation (3). Note that in this case, the actual action space is the one formed by the feature vectors $\Phi(a)$ for $a\in\mathcal{A}_{\operatorname*{full}}$.
4.1 Varying-dependence Actions
Assuming the outcome function is generated from a kernel inherently implies that the expected outcomes of actions are correlated. To study the effect of varying correlation levels among actions, we conduct the following experiments:
Sampling Outcome Functions from a Kernel.
We use the stationary RBF kernel (Williams and Rasmussen, 2006), and provide additional illustrations using the non-stationary Gibbs kernel in Supplementary:
[…]
where $l$ is a length-scale parameter that control the dependence between actions. By (Kanagawa et al., 2018, Theorem 4.12), we sample functions from a RKHS by first constructing the kernel matrix $\mathbf{K}$ with entries $\mathbf{K}_{a,a^{\prime}}=k(a,a^{\prime})$ for $a,a^{\prime}\in\mathcal{A}_{\operatorname*{full}}$, and then drawing $f\sim\mathcal{N}(0,\mathbf{K})$.
Figure 1: Performance comparison on stochastic combinatorial optimization in Equation (6), selecting $K=5$ distinct actions from 15 grid points, with outcome functions $f(a):=\mu_{a}$ sampled from an RBF kernel at varying length-scales. Left: Expected regret for our method (green), Thompson Sampling (yellow), UCB (orange), and Successive Halving (SH, blue), averaged over 50 repetitions at each length-scale. TS/UCB are run for 3000 rounds; SH is given a budget of 37,000 pulls. Right: SH’s expected regret (blue) and pull counts (yellow bars) per round. Our method’s performance (green) is shown for comparison. SH requires nearly 10,000 pulls—about 3× the number of super arms—to match our method, making it infeasible for large $\mathcal{A}_{\operatorname*{full}}$.
Returning to the objective stated in the Introduction, we aim to find a subset $\mathcal{A}$ of cardinality $K$ that minimizes the expected regret defined in Equation (1). Since the term $\mathbb{E}[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}]$ is a constant independent of $\mathcal{A}$, this is equivalent to the following optimization problem:
In general, this problem is non-trivial because the objective function involves an expectation over a collection of random variables whose distribution is neither specified nor directly accessible. Moreover, the expression inside the expectation is the maximum over a subset of potentially correlated random variables. As a result, small changes in $\mathcal{A}$ can lead to non-smooth variations in the maximum.
Our method is computationally efficient, avoiding the combinatorial complexity of the optimization problem in Equation (6). To assess effectiveness in identifying near-optimal solutions, we compare our method against Thompson Sampling (TS) and Upper Confidence Bound (UCB) in a simple example, using Successive Halving (SH) as a reference for approximating the optimal subset—though SH becomes impractical when the full action space is large.
We consider a fixed action space consisting of 15 grid points in [0,2][0,2][ 0 , 2 ], using an RBF kernel in Equation (5) to sample outcome functions while varying the length-scale $l\in\{0.5,1,\dots,4\}$. In our method, we run Algorithm 1 until $K=5$ distinct actions are selected, where the payoff of each action in a round is given by $\mu_{a}$. We use exhaustive search over the action space to find the optimal action $\arg\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}$. For other methods, we treat each $K$-tuple of actions as a super arm, and the payoff of each super arm in a round $\mathcal{A}$ is given by $\max_{a\in\mathcal{A}}\mu_{a}$. Thus, these methods are tailored to find the best super arm that maximizes $\mathbb{E}\left[\max_{a\in\mathcal{A}}\mu_{a}\right]$, aligning with the same objective. In TS/UCB methods, we adopt a bandit feedback setting: at each round, the decision-maker selects a super arm $\mathcal{A}$, observes the payoff $\max_{a\in\mathcal{A}}\mu_{a}$, and updates its policy accordingly, repeating this process for 3000 rounds, chosen to roughly match the number of super arms ($N=3003$). Since the payoff $\max_{a\in\mathcal{A}}\mu_{a}$ is unbounded, we assume Gaussian payoffs for both TS and UCB. The algorithms are summarized in Supplementary, with the prior of TS set to $\mathcal{N}(0,1)$. In the SH method, all super arms are evaluated using a fixed budget of arm pulls (we use the minimum budget required in Karnin et al. (2013), $N\log_{2}N\approx 37{,}000$) over a few rounds. In each round, the number of remaining super arms is halved. This process continues until only the best-performing super arm remains or the budget is exhausted. We evaluate expected regret of all methods over $10^{5}$ randomly sampled outcome functions.
The left subplot of Figure 1 reports the expected regret of our method (green), TS (yellow), UCB (orange) and SH (blue), where solid curves and error shades indicate the mean $\pm$ one standard deviation of expected regret over 50 repetitions. As shown in this subplot, the expected regret decreases as the length-scale increases. This is because the length-scale $l$ of the RBF kernel controls the number of approximately independent actions in the process $\{\mu_{a}\}_{a\in\mathcal{A}_{\operatorname*{full}}}$: When actions $a$ and $a^{\prime}$ are far apart relative to the length-scale, the kernel value becomes very small, so the covariance between the function values $\mu_{a}$ and $\mu_{a^{\prime}}$ is close to zero.
Regarding computational complexity: in a given bandit instance $\theta$, identifying the optimal action $a^{*}(\theta):=\arg\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}(\theta)$ is a well-studied problem in both best-arm identification and GP optimization. However, if computing $a^{*}(\theta)$ is not computationally efficient, then identifying the optimal super arm, as in SH—whose search space grows combinatorially—is infeasible. Next, in the right subplot of Figure 1, we examine sample complexity. We run our method (green) and Successive Halving (SH, blue) for 50 additional repetitions, using a fixed length-scale $l=1$. Since SH proceeds in rounds, we track the current best super arm and the total number of super arm pulls so far after each round. The expected regret of the current best super arm, is evaluated over another $10^{5}$ randomly sampled outcome functions. The blue solid curves and shades for SH represent the mean $\pm$ one standard deviation of expected regret over 50 repetitions over each SH round. The yellow bars (corresponding to the right y-axis) show the number of super arm pulls by SH at the end of each round. For comparison, the green curve and shaded region represent the mean $\pm$ one standard deviation of expected regret for our method over the same 50 repetitions. As shown, SH’s performance roughly matches ours after three rounds, requiring 9005 super arm pulls—approximately three times the number of super arms. If the action space $|\mathcal{A}_{\operatorname*{full}}|$ is already large, running SH is intractable.
4.2 Conclusion and Future Work
We proposed a framework for selecting a subset of correlated actions, modeling payoff correlations with a Gaussian process and extending to the sub-Gaussian case. A simple algorithm was introduced and shown to effectively identify near-optimal subsets. A key direction for future work is developing a stopping criterion. When the subset size is flexible but sampling new bandit instances is costly, ideas from species discovery Roswell et al. (2021) may be useful—treating the discovery of a new action as analogous to discovering a new species. One could use Turing’s formula to estimate the probability of finding a new action, or frame the process as a one-armed bandit that stops when the discovery probability becomes sufficiently low.
Appendix A Connection to Bayesian Regret of Multi-Armed Bandits
Given a subset $\mathcal{A}\subseteq\mathcal{A}_{\operatorname*{full}}$, let $\mu^{*}(\mathcal{A},\theta):=\max_{a\in\mathcal{A}}\mu_{a}(\theta)$ be the highest expected outcome in a bandit instance $\theta$. We can equivalently rewrite the regret in Equation (1):
$\textnormal{Regret}(\theta)=\mu^{*}(\mathcal{A}_{\operatorname*{full}},\theta)-\mu^{*}(\mathcal{A},\theta).$
Bandit Regret: In the classic multi-armed bandit setting [Auer et al., 2002a, Lattimore and Szepesvári, 2020], outcome of an arm $a\in\mathcal{A}_{\operatorname*{full}}$ in a bandit instance $\theta$ is generated i.i.d. according to an unknown but fixed law with expectation $\mu_{a}(\theta)$. A decision-maker would follow a policy $\pi$ that chooses the next arm based on the sequence of rounds and obtained outcomes. The bandit regret of policy $\pi$ in this bandit environment $\theta$ over $N$ rounds is defined as the expected difference in cumulative outcomes over $N$ rounds between always selecting the optimal action and following a policy $\pi$ to choose actions. Bayesian bandit regret [Agrawal and Goyal, 2012, Lattimore and Szepesvári, 2020] is the expectation of this bandit regret, taken over the randomness in the distribution of bandit instances, and the action chosen by the policy:
@@ Appendix A Connection to Bayesian Regret of Multi-Armed Bandits
$\begin{split}\textnormal{BanditRegret}(\theta):=&\mu^{*}(\mathcal{A}_{\operatorname*{full}},\theta)-\sum_{t=1}^{N}\mu_{A_{t}}(\theta),\\ \textnormal{BayesianBanditRegret}:=&\mathbb{E}\!\left[\mu^{*}(\mathcal{A}_{\operatorname*{full}},\theta)-\sum_{t=1}^{N}\mu_{A_{t}}(\theta)\right],\end{split}$
where $A_{t}$ is the action chosen by policy $\pi$ in round $t$. Since we consider settings where only a subset of the full action space is accessible, the action chosen by the policy can only from the subset $\mathcal{A}\subseteq\mathcal{A}_{\operatorname*{full}}$. This quantifies the expected difference in cumulative outcomes over $N$ rounds between always selecting the optimal action from the full action set $\mathcal{A}_{\operatorname*{full}}$ and following policy $\pi$ to choose actions from the restricted subset $\mathcal{A}$.
Next, we establish the connection between the regret defined in this paper and the Bayesian bandit regret.
A.1 Optimal policy
If $\pi$ is the optimal policy, it selects the best action in $\mathcal{A}$ in every round. Therefore, the regret of only having access to $\mathcal{A}$ instead of $\mathcal{A}_{\operatorname*{full}}$ for $N$ rounds, is determined by the optimal actions in $\mathcal{A}$ and $\mathcal{A}_{\operatorname*{full}}$:
[…]
A.2 Thompson sampling policy
If $\pi$ is the Thompson sampling policy, the Bayesian regret is bounded by
[…]
$=$ $\mathbb{E}\left[\sum_{t=1}^{N}\left(\mu^{*}(\mathcal{A})-\mu_{A_{t}}+\mu^{*}(\mathcal{A}_{\operatorname*{full}})-\mu^{*}(\mathcal{A})\right)\right]$
$\leq$ $C\sqrt{|\mathcal{A}|\cdot N\log N}+N\cdot\mathbb{E}_{\theta}[\textnormal{Regret}],$
where $|\mathcal{A}|$ denotes the cardinality, and $C>0$ is a constant. The inequality uses the Bayesian (bandit) regret of Thompson sampling (Theorem 36.1 in Lattimore and Szepesvári [2020]). Note that with access to the full action space, the Bayesian regret is bounded by $C\sqrt{|\mathcal{A}_{\operatorname*{full}}|\cdot N\log N}$. However, this can be worse than restricting to a subset $\mathcal{A}$ when $\mathbb{E}_{\theta}[\textnormal{Regret}]$ is sufficiently small and the time horizon $N$ is finite. Intuitively speaking, using a well-chosen subset of representative actions instead of the full action space can reduce Bayesian regret in multi-armed bandits.
Appendix B Proof of Lemma 4
@@
where the equality is the probability of missing the cluster $r$ for $K$ times, the first inequality uses the definition of $R_{\epsilon}$. The last inequality uses $\log(1-\epsilon)^{K}=K\log(1-\epsilon)$ and the inequality $\log(1-\epsilon)\leq-\epsilon$ for $\epsilon<1$ [TOPSØE1, 2007]. Therefore,
@@
where the first inequality uses union bound, the second inequality uses Equation (7). The last inequality uses the fact that there could be at most $\lfloor 1/\epsilon\rfloor$ clusters in $R_{\epsilon}$. ∎
Appendix C Proof of Theorem 7
@@
Consider a partition :=$\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space and an arbitrary reference subset $\mathcal{A}$. For each $\ell\leq m$, define a Gaussian process $\left\{Z_{a}\right\}_{a\in r_{\ell}}$ where $Z_{a}:=\mu_{a}-\mu_{a_{\ell}}$. Define a non-negative random variable
@@
[…]
@@
where the first equality uses integrated tail formula of expectation (cf. Lemma 1.6.1 of Vershynin [2018]). The last equality set $u_{0}:=\epsilon\sqrt{\log c}$. ∎
Lemma 15 (Borell-TIS inequality; Lemma 2.4.7 of [Talagrand, 2014]).
@@
Consider a partition :=$\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space, with $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)$, and an arbitrary reference subset $\mathcal{A}$. Then, for some constant $C>0$,
@@
where the first equality follows from the definition of regret in Equation (1), and the second equality follows from the assumption that $a^{*}(\theta)\in r_{\ell}$. The inequality holds because $a_{\ell}\in\mathcal{A}$, and hence $\max_{a\in\mathcal{A}}\mu_{a}\geq\mu_{a_{\ell}}$. The final equality follows from the definition of $Y_{\ell}$. On the other hand, according to Lemma 13, under Assumption 6, $\operatorname*{Regret}=Y_{\ell}$, and thus $\operatorname*{Regret}\geq\min_{\ell^{\prime}\leq m}Y_{\ell^{\prime}}$, where the cluster $r_{\ell}$ contains the optimal action.
@@
where the first inequality uses Equation (8). The second inequality uses $\max_{\ell\leq m}Y_{\ell}\leq\max_{\ell\leq m}\mathbb{E}Y_{\ell}+\max_{\ell\leq m}\left|Y_{\ell}-\mathbb{E}Y_{\ell}\right|$, since $Y_{\ell}\leq\mathbb{E}Y_{\ell}+\left|Y_{\ell}-\mathbb{E}Y_{\ell}\right|$. The last inequality uses Equation (9) and the identity $\mathbb{E}[Y_{\ell}]=\mathbb{E}[\max_{a\in r_{\ell}}\mu_{a}]$.
@@
where the first inequality uses Equation (8) under Assumption 6. The second inequality uses $\min_{\ell\leq m}Y_{\ell}\geq\min_{\ell\leq m}\mathbb{E}Y_{\ell}-\max_{\ell\leq m}\left|Y_{\ell}-\mathbb{E}Y_{\ell}\right|$, since $Y_{\ell}\geq\mathbb{E}Y_{\ell}-\left|Y_{\ell}-\mathbb{E}Y_{\ell}\right|$. The last inequality uses Equation (9) and the identity $\mathbb{E}[Y_{\ell}]=\mathbb{E}[\max_{a\in r_{\ell}}\mu_{a}]$. ∎
Appendix D Proof of Theorem 8

### v3

@@ Appendix B Extra Regret Bounds of Algorithm 1
$\begin{split}\mathbb{E}_{\theta,\mathcal{A}}[\operatorname*{Regret}]\geq\left(\min_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]-C\epsilon\sqrt{\log|\mathcal{R}|}\right)\times c\cdot\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\right)^{1/2}.\end{split}$
@@
Let $\gamma_{2}(\mathcal{S})$ denote Talagrand’s chaining functional [34] for a set $\mathcal{S}\subset\mathbb{R}^{n}$, $n\in\mathbb{N}$, equipped with the Euclidean norm. It is the sharpest known bound that for a sub-Gaussian process indexed by $\mathcal{S}$, there exists a constant $c>0$ such that $\mathbb{E}\left[\max_{a\in\mathcal{S}}\mu_{a}\right]\leq c\gamma_{2}(\mathcal{S})$. Using this functional, we establish a regret bound for our algorithm:
@@
Let $\{\mu_{a}\}_{a\in\mathcal{A}_{\operatorname*{full}}}$ be a mean-zero sub-Gaussian process. Consider a partition $\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space. Let $\mathcal{A}$ be output of Algorithm 1. Then, for a constant $C>0$,
@@
$\begin{split}\mathbb{E}_{\theta,\mathcal{A}}[\textnormal{Regret}]\leq C\sqrt{\log m}\cdot\max_{\ell\leq m}\gamma_{2}(r_{\ell})+\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu^{2}_{a}\right]\right)^{1/2}.\end{split}$
Proofs are in Appendices G and H, respectively.
Appendix C Proof of Lemma 4
@@
where the equality is the probability of missing the cluster $r$ for $K$ times, the first inequality uses the definition of $R_{\epsilon}$. The last inequality uses $\log(1-\epsilon)^{K}=K\log(1-\epsilon)$ and the inequality $\log(1-\epsilon)\leq-\epsilon$ for $\epsilon<1$ [35]. Therefore,
@@
where the first inequality uses union bound, the second inequality uses equation 6. The last inequality uses the fact that there could be at most $\lfloor 1/\epsilon\rfloor$ clusters in $R_{\epsilon}$. ∎
Appendix D Proof of Theorem 6
Under Assumption 5, the regret can be decomposed as follows; note that this Lemma is used only to derive lower bounds.
@@ Appendix D Proof of Theorem 6
Consider a partition $\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space and an arbitrary reference subset $\mathcal{A}$. For each $\ell\leq m$, define a Gaussian process $\left\{Z_{a}\right\}_{a\in r_{\ell}}$ where $Z_{a}:=\mu_{a}-\mu_{a_{\ell}}$. Define a non-negative random variable
@@
$\mathbb{E}X=$ $\int_{0}^{\infty}\Pr[X\geq u]\;du$
$=$ $\int_{0}^{u_{0}}\Pr[X\geq u]\;du+\int_{u_{0}}^{\infty}\Pr[X\geq u]\;du$
$\leq$ $u_{0}+\frac{1}{u_{0}}\int_{u_{0}}^{\infty}u\cdot\Pr[X\geq u]\;du$
$\leq$ $u_{0}+\frac{c}{u_{0}}\int_{u_{0}}^{\infty}u\cdot\exp{\left(-\frac{u^{2}}{\epsilon^{2}}\right)}\;du$
@@
where the first equality uses integrated tail formula of expectation (cf. Lemma 1.6.1 of [37]). The last equality set $u_{0}:=\epsilon\sqrt{\log c}$. ∎
Lemma 14 (Borell-TIS inequality; Lemma 2.4.7 of [34]).
@@
Consider a partition $\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space, with $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)$, and an arbitrary reference subset $\mathcal{A}$. Then, for some constant $C>0$,
@@
where the first equality follows from the definition of regret in equation 1, and the second equality follows from the assumption that $a^{*}(\theta)\in r_{\ell}$. The inequality holds because $a_{\ell}\in\mathcal{A}$, and hence $\max_{a\in\mathcal{A}}\mu_{a}\geq\mu_{a_{\ell}}$. The final equality follows from the definition of $Y_{\ell}$. On the other hand, according to Lemma 12, under Assumption 5, $\operatorname*{Regret}=Y_{\ell}$, and thus $\operatorname*{Regret}\geq\min_{\ell^{\prime}\leq m}Y_{\ell^{\prime}}$, where the cluster $r_{\ell}$ contains the optimal action.
@@
where the first inequality uses equation 7. The second inequality uses $\max_{\ell\leq m}Y_{\ell}\leq\max_{\ell\leq m}\mathbb{E}Y_{\ell}+\max_{\ell\leq m}\left|Y_{\ell}-\mathbb{E}Y_{\ell}\right|$, since $Y_{\ell}\leq\mathbb{E}Y_{\ell}+\left|Y_{\ell}-\mathbb{E}Y_{\ell}\right|$. The last inequality uses equation 8 and the identity $\mathbb{E}[Y_{\ell}]=\mathbb{E}[\max_{a\in r_{\ell}}\mu_{a}]$.
@@
where the first inequality uses equation 7 under Assumption 5. The second inequality uses $\min_{\ell\leq m}Y_{\ell}\geq\min_{\ell\leq m}\mathbb{E}Y_{\ell}-\max_{\ell\leq m}\left|Y_{\ell}-\mathbb{E}Y_{\ell}\right|$, since $Y_{\ell}\geq\mathbb{E}Y_{\ell}-\left|Y_{\ell}-\mathbb{E}Y_{\ell}\right|$. The last inequality uses equation 8 and the identity $\mathbb{E}[Y_{\ell}]=\mathbb{E}[\max_{a\in r_{\ell}}\mu_{a}]$. ∎
Appendix E Proof of Theorem 7
An important consequence of assuming $\mathbb{E}_{\theta}\left[\mu_{a}\right]=0$ for all $a\in\mathcal{A}_{\operatorname*{full}}$ is the following property: